# Entanglement in quenched extended Su-Schrieffer-Heeger model with anomalous dynamical quantum phase transitions

### v1

@@ III.0.2 EE in the extended SSH model
We begin the discussion of the extended SSH model by first considering the EE in its equilibrium phase diagram shown in Fig. 7. Recall that EE measures the information lack, as such it enables us to quantify the degree of non-locality for each phase. The figure suggests that the ground states of the $W_{2}$ and $W_{-1}$ phases are both more non-local than that of the $W_{0}$ and $W_{1}$ phases since most regions in $W_{2}$ and $W_{-1}$ phases have higher or approximately equal entropy values than those of $W_{0}$ and $W_{1}$ phases, regardless of the subsystems we choose. Therefore, it is natural to separate $W_{2}$ and $W_{-1}$ as one class of phase, while the other two being another class of phase. Another reason why we propose these two different regions is due to symmetry. If one performs the following mapping $(t_{c},t_{d},L_{A}=\alpha)\mapsto(t_{d},t_{c},L_{B}=\alpha)$, where $\alpha\in\mathbb{N}$, the equilibrium EEs remain unchanged. In another perspective, the EEs for $L_{A}=\alpha$ and $L_{B}=\alpha$ differ only by a reflection along the $t_{d}=t_{c}$ diagonal in Fig. 7. Therefore, we suggest categorizing $W_{0},W_{1}$ as class-1 phases, and $W_{2},W_{-1}$ as class-2 phases, where phases from class 2 have higher degree of non-locality than those in class 1. Phases in the same class have equal degree of non-locality. Such a classification of the phases is necessary to extract the generic features of EE around DQPT. This may serve as an alternative approach to probe entanglement physics for DQPTs with $n_{k^{*}}>1$ which was previously found with no apparent connection between EE and DQPTs via conventional approach [54]. We will demonstrate quenches between or within classes have distinctive EE properties near DQPT in the following discussion.
Figure 7: Entanglement entropy per unit cell with various subsystems (a) $L_{A}=1$, (b) $L_{B}=1$, (c) $L_{A}=2$, and (d) $L_{B}=2$ in the equilibrium extended SSH model. The top color bars are meant for (a) and (b), while the bottom ones are meant for (c) and (d). The simulation was performed with $L=120$.
@@
Class 1 $|\Delta W|=n_{k^{*}}$: quench dependent $|\Delta W|=n_{k^{*}}$: max for deep quench
$|\Delta W|<n_{k^{*}}$: max/min
Class 2 $|\Delta W|=n_{k^{*}}$: always min $|\Delta W|=n_{k^{*}}$: max for deep quench
$|\Delta W|<n_{k^{*}}$: always min
@@
The first diagonal block in Table 2 corresponds to quenches within phases from class 1. We notice that the behavior of EE in the vicinity of DQPT is mostly similar to that of the original SSH model, but with certain additional features. The quenches involving the $3k^{*}$ anomalous region, i.e. $\Delta W<n_{k^{*}}$, have similar features as the original quenched SSH model. For example, Fig. 8 shows the time evolution of EE for $L_{A}=1$ along some selected quench paths from $W_{1}$ to $W_{0}$ phases, enclosing a dynamical phase boundary where $n_{k^{*}}$ changes within the same equilibrium phase. The EE in the $3k^{*}$ region exhibits a local minimum near DQPT for $L_{A}=1$ (comparing Fig. 6(a) and 8(b)).
@@
For the $1k^{*}$ regular region, i.e. $\Delta W=n_{k^{*}}$, the behavior of the entanglement around DQPT is, in general, quench dependent. The EE may be a minimum, a maximum, or neither of them around DQPT. The inconsistency is due to the fact that the quench is near the equilibrium phase boundary in the $1k^{*}$ regime, as compared to the $3k^{*}$ regime where the quench is well away from the $t_{d}=t_{c}$ phase boundary. Quenches to places near the equilibrium phase boundary sometimes may not yield consistent results due to the rapid change of properties of the systems. These observations may suggest that the introduction of $t_{c}$ and $t_{d}$ hoppings and equalizing the weight between $t_{a}$ and $t_{b}$ hopping have the effect of recovering the EE property near DQPT of the original SSH model only in the $3k^{*}$ region, but partially in the $1k^{*}$ region.
@@
There is an interesting feature regarding quenches to regular and anomalous region in the extended SSH model. Figure 8(b) shows the DQPT occurs at the second local minimum (sometimes the third or fourth and so on) of the EE for the $1k^{*}$ region, whereas for the $3k^{*}$ region, DQPT consistently occurs at the first local minimum of EE. In fact, we always find the DQPT in the regular region occurs at a later time, while DQPTs in anomalous region happen earlier. This echoes with Table 1 that regular regions often have longer $t^{*}$, while anomalous region has shorter ones. This often results in a finite gap in the critical time when one quenches along a path from the regular region to the anomalous region, for example the cases shown in Fig. 8. One may interpret the first minimum in the EE in the $1k^{*}$ region as a pre-development of a DQPT. Once the quench enters the $3k^{*}$ phase, DQPT emerges at the first minimum immediately.
@@
Figure 8: (a) Loschmidt rate functions and (b) time evolution of the EEs with subsystem $L_{A}=1$ for quenches $W_{1}\rightarrow W_{0}$ of the extended SSH model. The initial parameter is $(t_{c}^{i},t_{d}^{i})=(-0.5,1)$, quenching to the various final parameters $(t_{c}^{f},t_{d}^{f})$ that lies on the path $t_{d}^{f}=-t_{c}^{f}+0.1$. The exact value of final parameters are shown in the legend. Colored dashed lines indicate the first critical time for the corresponding quench cases. Here $L=80$. Initial and final points are indicated as green dot and cross respectively in the inset color map of (a).
@@
Figure 9: Time evolution of the entanglement entropies per unit cell with subsystem $L_{A}=2$ involving quenches between the two classes: (a) from class 2 to class 1 and (b) from class 1 to class 2. The initial point in (a) and (b) are $(t_{c}^{i},t_{d}^{i})=(-0.5,-2)$ and $(t_{c}^{i},t_{d}^{i})=(-0.5,1)$ respectively. Legends denote the final point $(t_{c}^{f},t_{d}^{f})$ of the quenches, tracing a path of (a) $t_{c}^{f}=-t_{d}^{f}$ and (b) $t_{c}^{f}=1$. The vertical dashed lines pin the first critical time for each of the quench cases. Here $L=80$. Initial and final points of quenches in (a) are indicated as green dots and crosses respectively in the inset panel.
@@
Figure 10: Time evolution of the entanglement entropies per unit cells for quenches within class 2. (a) From $(t_{c}^{i},t_{d}^{i})=(-0.5,-2)$ along the path $t_{d}^{f}=-0.5$, (b) $(t_{c}^{i},t_{d}^{i})=(-0.5,-2)$ along the path $t_{d}^{f}=2$ and (c) $(t_{c}^{i},t_{d}^{i})=(-1.2,-1.8)$ along the path $t_{d}^{f}\approx-0.93t_{c}^{f}-1.16$. All subsystems are chosen to be $L_{A}=2$. The precise final quench parameter $(t_{c}^{f},t_{d}^{f})$ is indicated in the legends. Inset in (a) shows the critical times and the time when the EE for various subsystems achieves a maximum as a function of $t_{c}^{f}$. The colored dashed lines show the first critical time for the corresponding quench case. Here $L=80$. Inset color maps in (b) and (c) illustrate the quench cases, where final parameters are indicated as diamonds and crosses respectively while the initial points are indicated as green dots.
@@
We now discuss the result presented in the off-diagonal block in Table 2. Analogous to original SSH model, where $W_{0}$ and $W_{1}$ phase favors $t_{a}$ and $t_{b}$ hoppings respectively, we also need to determine the favorable hopping corresponding to $W_{2}$ and $W_{-1}$ phases before proceeding the discussion. Figure 2(a) suggests that $W_{2}$ and $W_{-1}$ phases are highly influenced by the $t_{d}$ and $t_{c}$ hopping terms respectively. Therefore, we propose $W_{2}$ and $W_{-1}$ phase favors $t_{d}$ and $t_{c}$ hoppings respectively.
@@
Figure 9(b) displays quenches from class 1 to class 2 (which corresponds to the (1,2) block in Table 2). On one hand, unfortunately, no analogy can be made from the protocol of inducing maximum/minimum near DQPT if one considers non-deep-phase quench. Deep-phase quench here refers to the situation where the postquenched Hamiltonian is dominated by one of the hoppings $t_{a}$, $t_{b}$, $t_{c}$ or $t_{d}$ over the rest. For example, an extremely large value of $t_{d}$ and small values of $t_{a}$, $t_{b}$ and $t_{c}$ is considered as deep inside $W_{2}$ phase. On the other hand, if one considers quenches involving the final parameter that is sufficiently deep inside the phase, it is possible to realize consistent maximum near DQPT for all the subsystems, even for the half-block case (not shown here).
@@
Figure 10 shows quenches between phases in class 2, which correspond to the second diagonal cell in Table 2. We first consider the case where $\Delta W\neq 0$. Similar to the case of class 1 quench to class 2, there is no general way to always attain an extremum if we use the protocol discussed concerning the original SSH model. The only consistent way to attain maximum is to quench sufficiently deep inside the phase. Figure 10(a) shows an example of deep-phase quench from $W_{2}$ to the $W_{-1}$ phase on the left. The inset shows the first critical time (the plus markers) and the time at which EE for various subsystem attains maxima (the hollow symbols) as a function of final $t_{c}^{f}$. It indicates the deeper we quench inside the phase (more negative $t_{c}^{f}$), the more the maximum EE and the critical time are aligned. Similar results are obtained for the quench from $W_{-1}$ to $W_{2}$ (not shown here). Note that the result holds for any subsystems including the case of a half-block (refer to Fig. 13(c), which shows the time evolution of the half-block correlation matrix spectrum for the quench $W_{2}\rightarrow W_{-1}$. The eigen-levels tend towards the Fermi level around DQPT results in higher EE. More detail will be discussed in the next section).
One can realize another type of quenches that stay within class-2 phases. Unlike all the quenches we mentioned above, this quench has $\Delta W=0<n_{k^{*}}$, which is a type-B anomalous DQPT. Figure 10(b) and 10(c) shows quenches between the two $W_{2}$ phases, with the former quench across the equilibrium phase boundary, while the latter does not. Using the protocol discussed in the original SSH model to choose the appropriate subsystem ($L_{A}=2$), these two varieties of quenches attain a local minimum in the vicinity of DQPT. Similar result can be observed if one changes the subsystem to $L_{B}=2$ for quenches within the $W_{-1}$ phase. Recall that the EE represents the information lack in a particular subsystem with respect to the whole system. The result suggests that in the vicinity of DQPT, the subsystem with 2 unit cells contain more information about the whole system as compared to $t=0$. A possible application is that one may be able to extract significant amount of the whole system’s information by just performing measurements on a small subsystem around DQPT of this type, without probing larger subsystems.
In Appendix A.1, the model is extended to include even longer range of hopping and the time evolution of the EE for some quench cases are presented. The results suggest the generalizability of our findings.
@@ IV Correlation Matrix Spectra
Figure 12: Time evolution of the correlation matrix spectra for quenches (a) $(-0.5,1)\rightarrow(1.01,-0.61)$, $W_{1}\rightarrow W_{0}$ inside $1k^{*}$ region; (b) $(-0.5,1)\rightarrow(1.58,-1.18)$, $W_{1}\rightarrow W_{0}$ inside $3k^{*}$ region; (c) $(-1.2,-1.8)\rightarrow(-0.5,-1.6)$, $W_{2}^{-}\rightarrow W_{2}^{-}$ without DQPT and (d) $(-1.2,-1.8)\rightarrow(-0.5,-0.9)$, $W_{2}^{-}\rightarrow W_{2}^{-}$ with DQPT for an $L=80$ system. Top, middle and bottom rows represent subsystems $L_{A}=2,4,40$ respectively. Colored dashed lines in (d) indicate first appearance of critical times. The first occurrence of the critical times for (a) and (b) are $t_{1}^{*}=1.1642$ and $(t_{1}^{*},t_{2}^{*},t_{3}^{*}=0.5333,0.5661,0.8962)$ respectively and are not shown in the plot.
@@
In the original SSH model, it has been found that the eigen-levels of the correlation matrix cross at the Fermi level if one quenches from the trivial $W_{0}$ phase to the topological $W_{1}$ phase and no crossings occur for quenches within the trivial phase [96]. Similar phenomenon happens for the trivial to topological phase quench in a model of the same Altland-Zirnbauer BDI class as the SSH model [48]. Moreover, persistent double degeneracy at the Fermi level is observed for quenches from topological to trivial phase [48]. The literature further proposed a dynamical quantity, termed entanglement echo, to indicate the crossings via a sudden jump in its rate function at the corresponding times. We found that the entanglement echo can behave very differently for different subsystem sizes and it does not always resemble the evolution of the Loschmidt rate in large subsystems, unlike the cases shown in Ref. [48] (see Appendix C).
@@
The above phenomena in the correlation matrix spectrum are not limited to the mentioned quench scenarios in the extended SSH model. Figure 11 shows the correlation matrix spectrum for quenches from each phase in the extended model. In addition to the consistent features found in the literature [96, 48] for $W_{1}\leftrightarrow W_{0}$ phase quenches (not presented here), we also observe similar features in $W_{-1}\leftrightarrow W_{0}$ phase quenches, where there are level crossings in the quench from the trivial to the topological phase (Fig. 11(a)), and a persistent double degeneracy at the $\xi_{j}(t)=0.5$, i.e. the Fermi level for the opposite quench (Fig. 11(d)). The latter can be deduced by noting that the Fermi level is doubly degenerated initially in the $W_{-1}$ phase and its evolution is a constant over time. Such a persistent degeneracy can also be observed for quenches between topological phases as shown in Fig. 11(b). Furthermore, we notice that level crossings occur when the system is quenched from the topological $W_{2}$ phase to the trivial phase, as shown in Fig. 11(c). Note that the Fermi level is four-fold degenerate initially in the $W_{2}$ phase and evolve into a pair of doubly degenerated levels until they cross at the Fermi level again. The crossing results in an instantaneous four-fold degeneracy of the Fermi level, retaining the number of edge modes in $W_{2}$ phases at the crossing times.
@@
We further investigate the correlation matrix spectrum’s evolution of different subsystem sizes for quenches starting from the $W_{1}$ and $W_{2}$ phases regarding the anomalous regions found in Fig. 2. Figure 12(a) and 12(b) present the correlation matrix spectrum for $W_{1}\rightarrow W_{0}$ quench in the regular and anomalous region respectively. Level crossing in the half-block’s correlation matrix spectrum is observed in $3k^{*}$ phase but not in $1k^{*}$ phase (bottom panels of Fig. 12(a) and 12(b)). The cases for $L_{A}=2$ behave the same but in a more drastic manner – level crossings barely occur when quenched to $1k^{*}$ region (top panel of Fig. 12(a)), whereas the near-Fermi levels frequently cross each other for quenches to $3k^{*}$ region, as shown in the top panel of Fig. 12(b). Subsystems $L_{A}=4$ seem to be insensitive to this kind of dynamical change, where there is no significant distinct feature in their correlation matrix spectra, as observed from the middle panels of Fig. 12(a) and 12(b). Overall, the quench to the anomalous regions triggers more frequent correlation redistributions for both very small and very large subsystems.
@@
On the contrary, one witnesses the rapid reorganization of correlation in a regular DQPT involving the class 2 $\rightarrow$ class 2 deep quench. Figure 13 shows the time evolution of the correlation matrix for a deep quench from $W_{2}^{-}$ to $W_{-1}^{-}$ phase. Note that for such kind of deep quenches, the originally three distinct critical times become close to each other due to the increasingly large energy each k mode possesses for $t_{c}^{f}\gg 1$ or $t_{d}^{f}\gg 1$. One observes level crossings around the critical times for all the subsystems concerned. Moreover, especially for $L_{A}=2,4$, the upper and lower levels also have their closest approach to the Fermi level around the first critical time. Same phenomenon can be observed for quenches to $W_{-1}^{+}$ phase. This is consistent with the observation that the entanglement entropy attains a local maximal around the first critical times as shown in Fig. 10(a). In the vicinity of DQPT, the system experiences a temporary redistribution of correlation in the subsystem, where a short instant of momentum uncertainty is introduced by the formation of quasiparticle modes at the Fermi level resulting in a restoration of the initial topological feature. This uncertainty in turn causes the loss of information to correctly describe the whole system and thus an increase in the entanglement entropy. The features regarding the correlation matrix spectrum presented in this section can also be observed in the extended SSH model with further neighbor hopping, as shown in Appendix A.2.

### v2

@@ III.0.2 EE in the extended SSH model
We begin the discussion of the extended SSH model by first considering the EE in its equilibrium phase diagram shown in Fig. 7. Recall that EE measures the information lack, as such it enables us to quantify the degree of non-locality for each phase. The figure suggests that the ground states of the $W_{2}$ and $W_{-1}$ phases are both more non-local than that of the $W_{0}$ and $W_{1}$ phases since most regions in $W_{2}$ and $W_{-1}$ phases have higher or approximately equal entropy values than those of $W_{0}$ and $W_{1}$ phases, regardless of the subsystems we choose. Therefore, it is natural to separate $W_{2}$ and $W_{-1}$ as one class of phase, while the other two being another class of phase. Another reason why we propose these two different classes is due to symmetry. If one performs the following mapping $(t_{c},t_{d},L_{A}=\alpha)\mapsto(t_{d},t_{c},L_{B}=\alpha)$, where $\alpha\in\mathbb{N}$, the equilibrium EEs remain unchanged. In another perspective, the EEs for $L_{A}=\alpha$ and $L_{B}=\alpha$ differ only by a reflection along the $t_{d}=t_{c}$ diagonal in Fig. 7. Therefore, we suggest categorizing $W_{0},W_{1}$ as class-1 phases, and $W_{2},W_{-1}$ as class-2 phases, where phases from class 2 have higher degree of non-locality than those in class 1. Phases in the same class have equal degree of non-locality. Such a classification of the phases is necessary to extract the generic features of EE around DQPT. This may serve as an alternative approach to probe entanglement physics for DQPTs with $n_{k^{*}}>1$ which was previously found with no apparent connection between EE and DQPTs via conventional approach [54]. We will demonstrate quenches between or within classes have distinctive EE properties near DQPT in the following discussion.
Figure 7: Entanglement entropy per unit cell with various subsystems (a) $L_{A}=1$, (b) $L_{B}=1$, (c) $L_{A}=2$, and (d) $L_{B}=2$ in the equilibrium extended SSH model. The top color bars are meant for (a) and (b), while the bottom ones are meant for (c) and (d). The black solid lines indicate the equilibrium phase boundary. The simulation was performed with $L=120$.
@@
Class 1 $n_{k^{*}}=|\Delta W|$ : quench dependent $n_{k^{*}}=|\Delta W|$ : max for deep quench
$n_{k^{*}}>|\Delta W|$: max/min
Class 2 $n_{k^{*}}=|\Delta W|$ : always min $n_{k^{*}}=|\Delta W|$ : max for deep quench
$n_{k^{*}}>|\Delta W|$: always min
@@
The first diagonal block in Table 2 corresponds to quenches within phases from class 1. We notice that the behavior of EE in the vicinity of DQPT is mostly similar to that of the original SSH model, but with certain additional features. The quenches involving the $3k^{*}$ anomalous region, i.e. $n_{k^{*}}>|\Delta W|$, have similar features as the original quenched SSH model. For example, Fig. 8 shows the time evolution of EE for $L_{A}=1$ along some selected quench paths from $W_{1}$ to $W_{0}$ phases, enclosing a dynamical phase boundary where $n_{k^{*}}$ changes within the same equilibrium phase. The EE in the $3k^{*}$ region exhibits a local minimum near DQPT for $L_{A}=1$ (see Fig. 8(b)).
@@
For the $1k^{*}$ regular region, i.e. $n_{k^{*}}>|\Delta W|$, the behavior of the entanglement around DQPT is, in general, quench dependent. The EE may be a minimum, a maximum, or neither of them around DQPT. The inconsistency is due to the fact that the quench is near the equilibrium phase boundary in the $1k^{*}$ regime, as compared to the $3k^{*}$ regime where the quench is well away from the $t_{d}=t_{c}$ phase boundary. Quenches to places near the equilibrium phase boundary sometimes may not yield consistent results due to the rapid change of properties of the systems. These observations may suggest that the introduction of $t_{c}$ and $t_{d}$ hoppings and equalizing the weight between $t_{a}$ and $t_{b}$ hopping have the effect of recovering the EE property near DQPT of the original SSH model only in the $3k^{*}$ region, but partially in the $1k^{*}$ region.
@@
There is an interesting feature regarding quenches to regular and anomalous region in the extended SSH model. Figure 8(b) shows the DQPT occurs at the second local minimum (sometimes the third or fourth and so on) of the EE for the $1k^{*}$ region, whereas for the $3k^{*}$ region, DQPT consistently occurs at the first local minimum of EE. In fact, we always find the DQPT in the regular region occurs at a later time, while DQPTs in anomalous region happen earlier. This echoes with Table 1 that regular regions often have longer $t^{*}$, while anomalous region has shorter ones. This often results in a finite gap in the critical time when one quenches along a path from the regular region to the anomalous region, for example the cases shown in Fig. 8. One may interpret the first minimum in the EE in the $1k^{*}$ region as a pre-development of a DQPT. Once the quench enters the $3k^{*}$ region, DQPT emerges at the first minimum immediately.
@@
Figure 8: (a) Loschmidt rate functions and (b) time evolution of the EEs with subsystem $L_{A}=1$ for quenches $W_{1}\rightarrow W_{0}$ of the extended SSH model. The initial parameter is $(t_{c}^{i},t_{d}^{i})=(-0.5,1)$, quenching to the various final parameters $(t_{c}^{f},t_{d}^{f})$ that lies on the path $t_{d}^{f}=-t_{c}^{f}+0.1$. The exact value of final parameters are shown in the legend. Colored dashed lines indicate the first critical time for the corresponding quench cases. Here $L=80$. Initial and final points are indicated as green dot and crosses respectively in the inset color map of (a).
@@
Figure 9: Time evolution of the entanglement entropies per unit cell with subsystem $L_{A}=2$ involving quenches between the two classes: (a) from class 2 to class 1 and (b) from class 1 to class 2. The initial point in (a) and (b) are $(t_{c}^{i},t_{d}^{i})=(-0.5,-2)$ and $(t_{c}^{i},t_{d}^{i})=(-0.5,1)$ respectively. Legends denote the final point $(t_{c}^{f},t_{d}^{f})$ of the quenches, tracing a path of (a) $t_{c}^{f}=-t_{d}^{f}$ and (b) $t_{c}^{f}=1$. The vertical dashed lines pin the first critical time for each of the quench cases. Here $L=80$. Initial and final points of quenches in (a) are indicated as green dot and crosses respectively in the inset panel.
@@
Figure 10: Time evolution of the entanglement entropies per unit cells for quenches within class 2. (a) From $(t_{c}^{i},t_{d}^{i})=(-0.5,-2)$ along the path $t_{d}^{f}=-0.5$, (b) $(t_{c}^{i},t_{d}^{i})=(-0.5,-2)$ along the path $t_{d}^{f}=2$ and (c) $(t_{c}^{i},t_{d}^{i})=(-1.2,-1.8)$ along the path $t_{d}^{f}\approx-0.93t_{c}^{f}-1.16$. All subsystems are chosen to be $L_{A}=2$. The precise final quench parameter $(t_{c}^{f},t_{d}^{f})$ is indicated in the legends. Inset in (a) shows the critical times and the times when the EE for various subsystems achieves a maximum as a function of $t_{c}^{f}$. The colored dashed lines show the first critical time for the corresponding quench case. Here $L=80$. Inset color maps in (b) and (c) illustrate the quench cases, where final parameters are indicated as diamonds and crosses respectively while the initial points are indicated as green dots.
@@
We now discuss the result presented in the off-diagonal block in Table 2. Analogous to original SSH model, where $W_{0}$ and $W_{1}$ phase favors $t_{a}$ and $t_{b}$ hoppings respectively, we also need to determine the favorable hopping corresponding to $W_{2}$ and $W_{-1}$ phases before proceeding the discussion. The equilibrium phase diagram in Fig. 7(a) suggests that $W_{2}$ and $W_{-1}$ phases are highly influenced by the $t_{d}$ and $t_{c}$ hopping terms respectively. Therefore, we propose $W_{2}$ and $W_{-1}$ phase favors $t_{d}$ and $t_{c}$ hoppings respectively.
@@
Figure 9(b) displays quenches from class 1 to class 2 (which corresponds to the (1,2) block in Table 2). On one hand, unfortunately, no analogy can be made from the protocol of inducing maximum/minimum near DQPT if one considers shallow phase quench. Deep-phase quench here refers to the situation where the postquenched Hamiltonian is dominated by one of the hoppings $t_{a}$, $t_{b}$, $t_{c}$ or $t_{d}$ over the rest. For example, an extremely large value of $t_{d}$ and small values of $t_{a}$, $t_{b}$ and $t_{c}$ is considered as deep inside $W_{2}$ phase. On the other hand, if one considers quenches involving the final parameter that is sufficiently deep inside the phase, it is possible to realize consistent EE maximum near DQPT for all the subsystems, even for the half-block case (not shown here).
@@
Figure 10 shows quenches between phases in class 2, which correspond to the second diagonal cell in Table 2. We first consider the case where $|\Delta W|\neq 0$. Similar to the case of class 1 quench to class 2, there is no general way to always attain an extremum if we use the protocol discussed concerning the original SSH model. The only consistent way to attain maximum is to quench sufficiently deep inside the phase. Figure 10(a) shows an example of deep-phase quench from $W_{2}$ to $W_{-1}$ phase on the left. The inset shows the first critical times (the plus markers) and the times at which EE for various subsystem attains maximum (the hollow symbols) as a function of $t_{c}^{f}$. It indicates the deeper we quench inside the phase (more negative $t_{c}^{f}$), the more the maximum EE and the critical time are aligned. Similar results are obtained for the quench from $W_{-1}$ to $W_{2}$ (not shown here). Note that the result holds for any subsystems including the case of a half-block (refer to the right panel of Fig. 13, which shows the time evolution of the half-block correlation matrix spectrum for the quench $W_{2}\rightarrow W_{-1}$. The eigen-levels tend towards the Fermi level around DQPT results in higher EE. More detail will be discussed in the next section).
One can realize another type of quenches that stay within class-2 phases. Unlike all the quenches we mentioned above, this quench has $n_{k^{*}}>\Delta W=0$, which is a type-B anomalous DQPT. Figures 10(b) and 10(c) show quenches between the two $W_{2}$ phases, with the former quench across the equilibrium phase boundary, while the latter does not. Using the protocol discussed in the original SSH model to choose the appropriate subsystem ($L_{A}=2$), these two varieties of quenches attain a local minimum in the vicinity of DQPT. Similar result can be observed if one changes the subsystem to $L_{B}=2$ for quenches within the $W_{-1}$ phase. Recall that the EE represents the information lack in a particular subsystem with respect to the whole system. The result suggests that in the vicinity of DQPT, the subsystem with 2 unit cells contain more information about the whole system as compared to $t=0$. A possible application is that one may be able to extract significant amount of the whole system’s information by just performing measurements on a small subsystem around DQPT of this type, without probing larger subsystems.
In Appendix A.1, the model is extended to include even longer range of hoppings and the time evolution of the EE for some quench cases are presented. The results suggest the generalizability of our findings.
@@ IV Correlation Matrix Spectra
Figure 12: Time evolution of the correlation matrix spectra for quenches (a) $(-0.5,1)\rightarrow(1.01,-0.61)$, $W_{1}\rightarrow W_{0}$ inside $1k^{*}$ region; (b) $(-0.5,1)\rightarrow(1.58,-1.18)$, $W_{1}\rightarrow W_{0}$ inside $3k^{*}$ region; (c) $(-1.2,-1.8)\rightarrow(-0.5,-1.6)$, $W_{2}^{-}\rightarrow W_{2}^{-}$ without DQPT and (d) $(-1.2,-1.8)\rightarrow(-0.5,-0.9)$, $W_{2}^{-}\rightarrow W_{2}^{-}$ with DQPT for an $L=80$ system. Top, middle and bottom rows represent subsystems $L_{A}=2,4,40$ respectively. Colored dashed lines in (d) indicate first appearance of critical times. The first occurrence of the critical times for (a) and (b) are $t_{1}^{*}=1.1642$ and $(t_{1}^{*},t_{2}^{*},t_{3}^{*})=(0.5333,0.5661,0.8962)$ respectively and are not shown in the plot.
@@
In the original SSH model, it has been found that the eigen-levels of the correlation matrix cross at the Fermi level ($\xi_{j}=0.5$) if one quenches from the trivial $W_{0}$ phase to the topological $W_{1}$ phase and no crossings occur for quenches within the trivial phase [96]. Similar phenomenon happens for the trivial to topological phase quench in a model of the same Altland-Zirnbauer BDI class as the SSH model [48]. Moreover, persistent double degeneracy at the Fermi level is observed for quenches from topological to trivial phase [48]. The literature further proposed a dynamical quantity, termed entanglement echo, to indicate the crossings via a sudden jump in its rate function at the corresponding times. We found that the entanglement echo can behave very differently for different subsystem sizes and it does not always resemble the evolution of the Loschmidt rate in large subsystems, unlike the cases shown in Ref. [48] (see Appendix C).
@@
The above phenomena in the correlation matrix spectrum are not limited to the mentioned quench scenarios in the extended SSH model. Figure 11 shows the correlation matrix spectrum for quenches from each phase in the extended model. In addition to the consistent features found in the literature [96, 48] for $W_{1}\leftrightarrow W_{0}$ phase quenches (not presented here), we also observe similar features in $W_{-1}\leftrightarrow W_{0}$ phase quenches, where there are level crossings in the quench from the trivial to the topological phase (Fig. 11(a)), and a persistent double degeneracy at $\xi_{j}(t)=0.5$, i.e. the Fermi level for the opposite quench (Fig. 11(d)). The latter can be deduced by noting that the Fermi level is doubly degenerated initially in the $W_{-1}$ phase and its evolution is a constant over time. Such a persistent degeneracy can also be observed for quenches between topological phases as shown in Fig. 11(b). Furthermore, we notice that level crossings occur when the system is quenched from the topological $W_{2}$ phase to the trivial phase, as shown in Fig. 11(c). Note that the Fermi level is four-fold degenerate initially in the $W_{2}$ phase and evolve into a pair of doubly degenerated levels until they cross at the Fermi level again. The crossing results in an instantaneous four-fold degeneracy of the Fermi level, retaining the number of edge modes in $W_{2}$ phases at the crossing times.
@@
We further investigate the correlation matrix spectrum’s evolution of different subsystem sizes for quenches starting from the $W_{1}$ and $W_{2}$ phases regarding the anomalous regions found in Fig. 2. Figures 12(a) and 12(b) present the correlation matrix spectrum for $W_{1}\rightarrow W_{0}$ quench in the regular and anomalous region respectively. Level crossing in the half-block’s correlation matrix spectrum is observed in $3k^{*}$ phase but not in $1k^{*}$ phase (bottom panels of Figs. 12(a) and 12(b)). The cases for $L_{A}=2$ behave the same but in a more drastic manner – level crossings barely occur when quenched to $1k^{*}$ region (top panel of Fig. 12(a)), whereas the near-Fermi levels frequently cross each other for quenches to $3k^{*}$ region, as shown in the top panel of Fig. 12(b). Subsystems $L_{A}=4$ seem to be insensitive to this kind of dynamical change, where there is no significant distinct feature in their correlation matrix spectra, as observed from the middle panels of Figs. 12(a) and 12(b). Overall, the quench to the anomalous regions triggers more frequent correlation redistributions for both very small and very large subsystems.
@@
On the contrary, one witnesses the rapid reorganization of correlation in a regular DQPT involving the class 2 $\rightarrow$ class 2 deep quench. Figure 13 shows the time evolution of the correlation matrix for a deep quench from $W_{2}^{-}$ to $W_{-1}^{-}$ phase. Note that for such kind of deep quenches, the originally three distinct critical times become close to each other due to the increasingly large energy each k mode possesses for $t_{c}^{f}\gg 1$ or $t_{d}^{f}\gg 1$. One observes level crossings around the critical times for all the subsystems concerned. Moreover, especially for $L_{A}=2,4$, the upper and lower levels also have their closest approach to the Fermi level around the first critical time. Same phenomenon can be observed for quenches to $W_{-1}^{+}$ phase. This is consistent with the observation that the entanglement entropy attains a local maximum around the first critical times as shown in Fig. 10(a). In the vicinity of DQPT, the system experiences a temporary redistribution of correlation in the subsystem, where a short instant of momentum uncertainty is introduced by the formation of quasiparticle modes at the Fermi level resulting in a restoration of the initial topological feature. This uncertainty in turn causes the loss of information to correctly describe the whole system and thus an increase in the entanglement entropy. The features regarding the correlation matrix spectrum presented in this section can also be observed in the extended SSH model with further neighbor hoppings, as shown in Appendix A.2.